OpenStudy (anonymous):

A logarithmic equation. I got it however , but i wish to share it with you all @mathslover :) Anyone could try to solve this problem

OpenStudy (anonymous):

\[\huge x ^{1+logx} = 10x\]

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

Anyone may solve this if they want . I got it

OpenStudy (anonymous):

I posted this this because some of them are doing logarithms and i am doing it too

OpenStudy (anonymous):

Its a tricky one. somewhat

OpenStudy (anonymous):

i thought this was for mathsolver :P

OpenStudy (anonymous):

Anyone here i specially tagged him because he is doing log

OpenStudy (anonymous):

ok ill give a direct answer :P

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

message me

OpenStudy (anonymous):

It is not so easy as it looks

mathslover (mathslover):

Its easy...

mathslover (mathslover):

@No.name - I think, the first step will be to take log both sides.

OpenStudy (anonymous):

yes

mathslover (mathslover):

Is the answer 0 ?

OpenStudy (anonymous):

zero dnt count :)

mathslover (mathslover):

O.O yeah.. I meant to say 1

mathslover (mathslover):

Sorry, i calculated log x = 0 So, x = 1

OpenStudy (anonymous):

haha no lol if x=1 then 1=10 O.o

mathslover (mathslover):

o.O leave it

mathslover (mathslover):

hmm, lol, is it 10

OpenStudy (anonymous):

i have solution but its irrational xD

mathslover (mathslover):

Hmm... I am getting 10 @No.name

OpenStudy (anonymous):

cool it work :O

mathslover (mathslover):

(1 + log x) (logx ) = log (10x) log x + (log x)^2 = log x + 1 (log x)^2 = 1 log x = +/- 1 log x can not be -1... So, log x = 1 thus, x = 10

OpenStudy (anonymous):

why its work mmm

OpenStudy (anonymous):

there are 2 answers

mathslover (mathslover):

One of them is 10?

OpenStudy (anonymous):

that's why its tricky

OpenStudy (anonymous):

yes one is 10

mathslover (mathslover):

hmm ... we might want to take the conditions here?

OpenStudy (anonymous):

nope its very straightforward

mathslover (mathslover):

Oh...

ganeshie8 (ganeshie8):

\(\large x^{\log_{10} 10 + \log_{10} x} = 10x\) \(\large x^{\log_{10} 10x} = x^{\log_{x} 10x}\) \(\large \log_{10} 10x = \log_{x} 10x\) \(\implies x = 10, ~0.1\)

OpenStudy (anonymous):

@ganeshie8 got it !

OpenStudy (anonymous):

One more question i have in log

mathslover (mathslover):

o.O yeah, I MUST not have avoided this : log (x ) = -1 and log(x) = 1 log(x) = -1 when ... x = 10^{-1} = 1/10

mathslover (mathslover):

OH.. :( I was almost there... but anyways, @ganeshie8 got it... great work and nice quest. @No.name :-)

mathslover (mathslover):

I went wrong at this pt. : "log x can not be -1..." it can be -1 :(

OpenStudy (anonymous):

yes

OpenStudy (anonymous):